Circular No. 7890

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 EJ_{57} (LINEAR) An apparently asteroidal object discovered by LINEAR was given the

designation 2002 EJ_{57} (observations on MPS 55777–55778). Additional observations of the object were located during routine processing at the Minor Planet Center, and these indicated that the orbit was unusual. The object was then placed on the NEO Confirmation Page, resulting in CCD observations by J. Tichá and M. Tichý at Kleť (diffuse with coma diameter ∼ 10′′ on May 1.8 UT) and by R. H. McNaught at Siding Spring (circular coma with diameter ∼ 7′′ on May 2.5) showing cometary activity.

2002 UT  α_{2000}  δ_{2000}  m_{2}
Mar. 13.20032  10^{h}24^{m}33.^{s}18  +8^{o}23′58′′.4  18.5

The available astrometry appears on MPEC 2002-J06, together with the following orbit that shows this to be a short-period comet:

T = 2001 Dec. 17.0076 TT  ω = 166.^{o}2200
e = 0.595362  Ω = 330.4054 }2000.0
q = 2.630672 AU  i = 4.9792
a = 6.501291 AU n^{o} = 0.0594572 P = 16.6 years

SUPERNOVA 2002cr IN NGC 5468 S. Nakano, Sumoto, Japan, reports the discovery by Reiki Kushida

(Yatsugatake South Base Observatory, 0.40-m Schmidt-Cassegrain f/5.7 reflector) of an apparent supernova on unfiltered CCD images taken on May 1.701 (at mag 16.5) and 2.467 UT (mag 16.2). SN 2002cr is located at α = 14^{h}06^{m}37^{s}.59, δ = −5^{o}26′21′′.9 (equinox 2000.0), which is 41′′ east and 50′′ north of the center of NGC 5468, in the faint northeast arm of the galaxy. Searches of images taken by Kushida during 1996–1997 (limiting mag 18.5) and other photographic atlases show nothing in this position.

COMET C/2002 E2 (SNYDER-MURAKAMI) Visual m_{1} estimates: Mar. 26.86 UT, 11.1 (A. Pearce, Nedlands, W.

Australia, 0.41-m reflector); Apr. 5.12, 10.5 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); 10.08, 10.7 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 13.06, 10.6 (A. Baransky, Kiev, Ukraine, 0.36-m reflector); 21.76, 11.0 (Y. Nagai, Yamanashi, Japan, 0.32-m reflector); 30.91, 11.1 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

2002 May 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green